OpenStudy (anonymous):

System of equation 12x+1y=300 15x+1y=4140

OpenStudy (anonymous):

subtract the one from the other.

OpenStudy (anonymous):

\(\begin{cases} \\ 15x+1y=4140 \\12x+1y=300 \hspace{25pt}- \end{cases} \\-----------\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it will leave you with only x and numbers, so you can solve x. once you know x, y should be no problem anymore.

OpenStudy (anonymous):

Why did you you reorder the problem

OpenStudy (anonymous):

because it's easier to subtract smaller numbers from bigger numbers, but you can do it the other way around as well.

OpenStudy (anonymous):

The answer would be negative

OpenStudy (anonymous):

yes, but you'll end up with the same answer in the end either way.

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

you got it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

alright, good job.